Solve the problem using a system of equations. At the end of 2001, Subway topped McDonalds as the largest restaurant chain in the U.S. Subway operated 148 more restaurant than McDonalds, and together the two chains had 26,346 restaurants. How many restaurants did each company operate?:
Answer by stanbon(57347)

Solve the problem using a system of equations. At the end of 2001,
Subway topped McDonalds as the largest restaurant chain in the U.S.
Subway operated 148 more restaurant than McDonalds, and together the two chains had 26,346 restaurants.
How many restaurants did each company operate?:
Equations:
m + s = 26346
s = m + 148
--------------------
Substitute to solve for "m":
m + (m+148) = 26346
2m = 26198
m = 13099 (# of outlets McDonals has)
m+148 = 13247 (# of outlets subway has)
===========================================
Cheers,
Stan H.